# How to Train Long-Context Language Models (Effectively)

### v1

@@ Abstract
We study continued training and supervised fine-tuning (SFT) of a language model (LM) to make effective use of long-context information. We first establish a reliable evaluation protocol to guide model development—instead of perplexity or simple needle-in-a-haystack (NIAH) tests, we use a broad set of long-context tasks, and we evaluate models after SFT with instruction data as this better reveals long-context abilities. Supported by our robust evaluations, we run thorough experiments to decide the data mix for continued pre-training, the instruction tuning dataset, and many other design choices. We find that (1) code repositories and books are excellent sources of long data, but it is crucial to combine them with high-quality short data; (2) training with a sequence length beyond the evaluation length boosts long-context performance; (3) for SFT, using only short instruction datasets yields strong performance on long-context tasks. Our final model, ProLong-8B, which is initialized from Llama-3 and trained on 40B tokens, demonstrates state-of-the-art long-context performance among similarly sized models at a length of 128K. ProLong outperforms Llama-3.1-8B-Instruct on the majority of long-context tasks despite having seen only 5% as many tokens during long-context training. Additionally, ProLong can effectively process up to 512K tokens, one of the longest context windows of publicly available LMs. [^1] Our code, data, and models are available at https://github.com/princeton-nlp/ProLong.
@@ 1 Introduction
While many works have focused on extending the context length of pre-trained LMs with minimal training (Chen et al., 2023; Peng et al., 2024), Fu et al. (2024) find that the above methods cannot even perform the simple needle-in-a-haystack (NIAH; Kamradt, 2024) task and it is necessary to continually train the LM on long documents for billions of tokens. Frontier open-source models, such as Llama-3.1 (Dubey et al., 2024) and Jamba (Lenz et al., 2024), also employ a long-context continued training stage, followed by supervised fine-tuning (SFT) on instruction data. We adopt the same setting and study continued training and SFT of a pre-trained LM for effective long-context use.
We first establish a reliable evaluation protocol to provide a meaningful signal for model development. Most existing works rely on either perplexity or NIAH for ablating training recipes. We demonstrate that neither is robust for guiding the development and opt for a broad range of downstream applications, such as retrieval-augmented generation (RAG), long-document summarization, and many-shot in-context learning (ICL). Importantly, we also conduct our evaluations after performing SFT, even for all our ablation runs on continued pre-training. We observe that, on some long-context tasks, performance gains only emerge after SFT, which means that best design choices can differ before and after SFT.
@@
2 Guide Model Development With Meaningful Evaluations
@@
2.1 Evaluate on diverse and realistic tasks
We first make the decision to use HELMET (Yen et al., 2024b) as our evaluation suite, as it is one of the most comprehensive long-context benchmarks, covering the following tasks:
@@ 2.1 Evaluate on diverse and realistic tasks
Overall, these diverse tasks reflect a range of long-context abilities including recall, reasoning, learning from context, and robustness to noisy inputs. Yen et al. (2024b) also show that HELMET produces model performance trends that are more consistent with human perceptions unlike other long-context benchmarks.
@@
We offer more details about the HELMET evaluation, including its careful choice of metrics, in §A.1. We did not use all tasks from HELMET for efficiency reasons and we also do not want to overfit to HELMET. If not otherwise specified, we average the performance for each category over all datasets and over evaluation lengths of 32K and 64K; for the final long-context score, we macro-average all categories.
@@
2.2 Evaluate after supervised fine-tuning
@@
We show the benchmarking results before and after SFT in Figure 2. Long-context evaluation shows clearer signals when it is conducted after SFT: (1) SFT shows that the model continues to improve with more training tokens on RAG and re-ranking, while the improvement is less clear or does not exist when evaluated before SFT. (2) SFT enables evaluation on realistic applications like QA and summarization, which require instruction following and have low performance before SFT. We also note that the variance from two random training runs is not substantially higher after the additional SFT phase. Therefore, unless otherwise specified, we report the long-context performance after SFT.
Figure 2: Improvements on RAG and re-ranking tasks are only observed when evaluating models after a supervised fine-tuning (SFT) phase on instruction data. The models are trained on the pre-training data mix by Fu et al. (2024). We report the mean and standard deviations over two training runs.
@@
2.3 Check that short-context performance is preserved
@@ 3.1 Code repositories and books are good sources of long-context data
Table 4: Impact of different long data sources, while keeping the 40% short data component fixed. Code repositories particularly helps recall, while books are more effective on re-ranking, ICL, and summarization. Mixing the two achieves the overall best performance.
@@ 3.2 Training only on long data hurts long-context performance
The ratio between short/long data is another crucial factor for downstream performance. Prior work either trains only on long data (Peng et al., 2024) or adds some short training data (Yen et al., 2024a; Fu et al., 2024). However, we are the first to systematically study the impact of short/long ratio.
@@ 4 Scaling the Size and Length of the Training Data
One might assume that we should train long-context models on the maximum sequence length that we want the model to support. Many works even emphasize extrapolation to even longer sequences at inference time (Press et al., 2022; Xiao et al., 2024b; a; Yen et al., 2024a; Chen et al., 2023). In contrast, we observe that training on a longer sequence length (512K tokens) substantially improves the long-context performance at a shorter evaluation length (64K tokens).
@@ 5 Supervised Fine-Tuning for Long-Context LMs
We first establish that UltraChat outperforms Tulu-v2 and ShareGPT in Table 22. We therefore use it when studying the ratio of synthetic long-context instruction data in Table 8. Surprisingly, we find that adding synthetic data does not improve the performance on these very long-context tasks, and adding even as little as 1% synthetic data hurts the performance in our setting. Therefore, we use only short-context UltraChat data for SFT of our final ProLong model.
@@
11% Tulu-v2, 11% StackExchange,
8% Wikipedia, 8% OpenWebMath, 8% ArXiv
@@ 6.2 ProLong performance
We present the final HELMET evaluation results of ProLong in Table 10. We compare to a number of frontier long-context LMs, namely MegaBeam [^6] https://huggingface.co/aws-prototyping/MegaBeam-Mistral-7B-512k., Llama-3.1 (Dubey et al., 2024), Qwen2 (Yang et al., 2024a), Phi-3 (Abdin et al., 2024), Mistral-Nemo [^7] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407., Jamba-1.5 (Lenz et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024), Gemini-1.5 (Reid et al., 2024), and GPT-4o (Achiam et al., 2023).
ProLong outperforms all 10B-scale models on our long-context evaluation. Notably, ProLong outperforms Llama-3.1-8B-Instruct on all categories except summarization. ProLong achieves this with only 5% of Llama-3.1’s long-context data budget (40B vs. 800B tokens). We also showcase the strength of ProLong with several QA examples in Table 23.
Table 10: Our main evaluation results on HELMET (Yen et al., 2024b; details in §A.1). All results are averaged over sequence lengths of 32K, 64K, and 128K. For all models, we use the corresponding instruction version. ProLong is one of the best performing 10B-scale LMs. The complete set of results can be found in §C.
Model Max Len. Recall RAG ICL Re-rank QA Summ. Avg.
ProLong (8B) 512K 99.4 66.0 81.1 33.2 40.8 40.5 60.2
MegaBeam-Mistral (7B) 512K 99.4 58.1 82.1 22.1 33.7 43.6 56.5
Meta-Llama-3.1 (8B) 128K 98.7 62.8 79.7 26.6 40.4 46.1 59.0
Qwen2 (7B) 128K 34.4 43.4 54.8 4.6 23.3 38.5 33.2
Phi-3-small (7B) 128K 74.8 60.6 82.0 18.5 34.1 42.4 52.1
Mistral-Nemo (12B) 128K 24.9 48.1 82.0 4.7 37.7 37.0 39.1
Jamba-1.5-Mini (12B/52B) 256K 87.7 61.3 88.4 25.9 42.0 38.6 57.3
Meta-Llama-3.1 (70B) 128K 98.5 65.9 80.0 39.4 47.2 51.1 63.7
Claude-3.5-Sonnet 200K 99.4 44.0 79.3 19.9 38.1 49.2 55.0
Gemini-1.5-Pro 2M 94.2 71.4 78.9 65.3 44.4 56.2 68.4
GPT-4o 128K 99.9 71.5 86.7 59.6 47.0 55.7 70.1
Since most existing models do not support more than 128K tokens, to showcase ProLong’s 512K context length, we stress test ProLong on the QA and summarization tasks from 32K to 512K888 In QA and summarization, we truncate the documents at the evaluation length; hence an effective long-context model should demonstrate better performance on longer lengths.. Table 11 shows that ProLong continues to improve at a longer context window.
@@ 7 Related Work
There have been many efforts in designing more efficient architectures, for example, linear attention/RNNs (Gu & Dao, 2023; Dao & Gu, 2024; Ma et al., 2022; Sun et al., 2023; Peng et al., 2023; Yang et al., 2024b), and alternative attention architectures (Rubin & Berant, 2023; Sun et al., 2024; Yen et al., 2024a). However, they often require training from scratch and many have the inherent limitations in terms of long-context recall (Jelassi et al., 2024; Arora et al., 2024). Recent works explore hybrid models (Waleffe et al., 2024; Lieber et al., 2024)) or distilling existing LMs into hybrid models (Wang et al., 2024) and show promising results.
@@
Many benchmarks have been proposed for long-context evaluation (Shaham et al., 2023; Hsieh et al., 2024; Krishna et al., 2023; Zhang et al., 2024a; An et al., 2024a; Bai et al., 2024b) There are works studying particular aspects of long-context LMs as well, such as positional bias (Liu et al., 2024b), in-context learning (Bertsch et al., 2024; Li et al., 2024b), and book-length summarization (Kim et al., 2024). In this work, we follow Yen et al. (2024b) for its diverse application coverage and reliable evaluations.
@@ 8 Conclusion
Although we try to ablate the major components of our training recipe, due to resource limitations, we cannot exhaust all aspects, such as the optimization hyperparameters and additional data mixtures. We also limit ourselves to the 10B-scale regime and the Llama-3 models, which may limit the generalizability of our findings and recipe. Another concern is that we are overfitting to the tasks chosen for model development—however, we do not directly train on those datasets and guiding model development with benchmark tasks has become a common practice in pre-trained LM development. We also show that our final recipe and model perform well on an additional evaluation dataset, NoCha.
@@ A.1 Evaluation
Table 13: The details for our long-context evaluation following HELMET (Yen et al., 2024b).
@@
Table 13 shows all the datasets we used for the long-context evaluation from HELMET (Yen et al., 2024b). Note that we did not use all the datasets from HELMET for efficiency reasons and we also do not want to overfit to HELMET. We highlight some of the evaluation protocol improvements that HELMET implemented compared to previous benchmarks here:
@@
  • Model-based evaluation for long-context QA and summarization. Instead of using traditional metrics like ROUGE (which has shown to be poorly indicative of the real model performance: Deutsch & Roth, 2021; Deutsch et al., 2022; Goyal et al., 2023; Chang et al., 2024), HELMET uses model-based evaluations to compare the reference answer and the model output. For QA, HELMET uses GPT-4o to score the model output given the question and the reference answer at a 0-3 scale. For summarization, HELMET takes a similar approach as Zhang & Bansal (2021); Gao et al. (2023): it first uses GPT-4o to decompose the reference summary into atomic claims; then it uses GPT-4o to check whether each reference atomic claim is covered by the model output (recall) and whether each sentence in the model output is covered by the reference summary (precision). Yen et al. (2024b) show that the model-based evaluation correlates with human perceptions significantly better than traditional metrics.
@@ A.3 Implementation details
We use various open-source packages and tools for the ProLong training and evaluation. We use PyTorch (Paszke et al., 2019) and Hugging Face transformers (Wolf et al., 2020) for the model training. We use mosaic-streaming (Mosaic ML, 2022) for loading and mixing the data and FlashAttention 2 (Dao, 2024) for efficient attention implementation. We implement sequence parallelism based on DeepSpeed-Ulysses (Jacobs et al., 2023) across groups of 8 GPUs on the same node. We only perform distributed attention if it is necessary, i.e., only on sequences of 512K length. For long-context evaluation, we use HELMET (Yen et al., 2024b) and for short-context evaluation, we use lm-eval-harness (Gao et al., 2021).

### v3

@@ Abstract
We study continued training and supervised fine-tuning (SFT) of a language model (LM) to make effective use of long-context information. We first establish a reliable evaluation protocol to guide model development—instead of perplexity or simple needle-in-a-haystack (NIAH) tests, we use a broad set of long-context downstream tasks, and we evaluate models after SFT as this better reveals long-context abilities. Supported by our robust evaluations, we run thorough experiments to decide the data mix for continued pre-training, the instruction tuning dataset, and many other design choices such as position extrapolation. We find that (1) code repositories and books are excellent sources of long data, but it is crucial to combine them with high-quality short-context data; (2) training with a sequence length beyond the evaluation length boosts long-context performance; (3) for SFT, using only short instruction datasets yields strong performance on long-context tasks. Our final model, ProLong-8B, which is initialized from Llama-3 and trained on 40B tokens, demonstrates state-of-the-art long-context performance among similarly sized models at a length of 128K. ProLong outperforms Llama-3.1-8B-Instruct on the majority of long-context tasks despite using only 5% as many tokens during long-context training. Additionally, ProLong can effectively process up to 512K tokens, one of the longest context windows of publicly available LMs. [^1] Our code, data, and models are available at https://github.com/princeton-nlp/ProLong.
@@ 1 Introduction
While many works have focused on extending the context length of pre-trained LMs with minimal training (Chen et al., 2023; Peng et al., 2024, inter alia), the above methods fail to solve even the simple needle-in-a-haystack (NIAH; Kamradt, 2024) task and it is necessary to continue LLM training on billions of tokens of long documents to learn this task robustly. Frontier open-source models, such as Llama-3.1 (Dubey et al., 2024) and Jamba (Lenz et al., 2024), also employ a long-context continued training stage, followed by supervised fine-tuning (SFT) on instruction data. We adopt the same setting and study continued training and SFT of a pre-trained LM for effective long-context use.
We first establish a reliable evaluation protocol to provide a meaningful signal for model development. Most existing works rely on either perplexity or NIAH for ablating training recipes. We demonstrate that neither is robust for guiding the development and opt for a broad range of downstream applications, such as retrieval-augmented generation (RAG), long-document summarization, and many-shot in-context learning (ICL). Importantly, we also conduct our evaluations after performing SFT, even for all our ablation runs on continued pre-training. We observe that, on some long-context tasks, performance gains only emerge after SFT, which means that best design choices can differ before and after SFT. We also check if the base model’s short-context performance is preserved.
@@
2 Guiding Model Development With Meaningful Evaluations
@@
2.1 Evaluating on diverse and realistic tasks
We first make the decision to use HELMET (Yen et al., 2025) as our evaluation suite, as it is one of the most comprehensive long-context benchmarks. For fast iteration, we only use a subset of HELMET tasks for model development:
@@ 2.1 Evaluating on diverse and realistic tasks
[…]
@@
We offer more details about the HELMET evaluation, including its careful choice of metrics, in §A.1. If not otherwise specified, we average the performance for each category over all datasets and over evaluation lengths of 32K and 64K; for the final long-context score, we macro-average all categories.
@@
2.2 Evaluating after supervised fine-tuning
@@
Figure 2: Some long-context improvements are only observed when evaluating after SFT. We report the mean and std over two training runs.
We show the benchmarking results before and after SFT in Figure 2. Long-context evaluation shows clearer signals when it is conducted after SFT: (1) SFT shows that the model continues to improve with more training tokens on RAG and re-ranking, while the improvement is less clear or does not exist when evaluated before SFT. (2) SFT enables evaluation on realistic applications like QA and summarization, which require instruction following and would otherwise fail completely. We also note that the variance from two random training runs is not substantially higher after the additional SFT phase. Therefore, unless otherwise specified, we report the long-context performance after SFT.
@@
2.3 Checking that short-context performance is preserved
@@ 3.1 Code repositories and books are good sources of long-context data
Table 4: Impact of different long data sources. Long-context performance is averaged over 32K and 64K lengths.
@@ 3.2 Training only on long data hurts long-context performance
The ratio between short/long data is another crucial factor for downstream performance. Prior work either trains only on long data (Peng et al., 2024) or adds some short training data (Yen et al., 2024; Fu et al., 2024). However, we are the first to systematically study the impact of short/long ratio.
@@ 3.3 Choosing a high-quality short-context mix is important
Comparison to prior efforts.
To confirm the effectiveness of our long-context data curation, we conduct a head-to-head comparison with the previous work by Fu et al. (2024) in §B.6. Our data mix significantly surpasses Fu et al. (2024) in both long-context and short-context tasks, underscoring the efficacy of ProLong.
@@ 4 Scaling the Size and Length of the Training Data
One might assume that we should train long-context models on the maximum sequence length that we want the model to support. Many works even emphasize extrapolation to even longer sequences at inference time (Press et al., 2022; Xiao et al., 2024b; a; Yen et al., 2024; Chen et al., 2023). In contrast, we observe that training on a longer sequence length (512K tokens) substantially improves the long-context performance at a shorter evaluation length (64K tokens).
@@
What is the benefit of training on longer sequences than used during evaluation? Here is our hypothesis: Assume that in order to solve a certain task (e.g., recall), a model has to be trained on examples of dependencies that span a distance of precisely $d$ tokens, i.e., there is no generalization between dependencies of different lengths. Also assume that these dependencies are equally likely to occur at any position in a sequence. Then, a document of length $nd$ will have $(n-1)(d-1)$ more dependencies of distance $d$ than $n$ documents of length $d$. While these assumptions do not hold in practice, this simplified model still provides intuition for our empirical findings.
@@ 5 Supervised Fine-Tuning for Long-Context LMs
We first establish that UltraChat outperforms Tulu-v2 and ShareGPT in Table 22. We therefore use it when studying the ratio of synthetic long-context instruction data in Table 8. Surprisingly, we find that adding synthetic data does not improve the performance on these very long-context tasks, and adding even as little as 1% synthetic data hurts the performance in our setting. We also verify that this phenomenon persists even when we use a more powerful data generator, such as Llama-3-70B (§B.5). Therefore, we use only short-context UltraChat data for SFT of our final ProLong model.
@@
11% Wikipedia, 11% StackExchange,
8% Tulu-v2, 8% OpenWebMath, 8% ArXiv
@@ 6.2 ProLong performance
We first verify that ProLong preserves the base model’s short-context performance in §B.7. We then present the final HELMET evaluation results of ProLong in Table 10. We use all available HELMET tasks here and please refer to Yen et al. (2025) for more details. We compare to a number of frontier long-context LMs, namely MegaBeam [^6] https://huggingface.co/aws-prototyping/MegaBeam-Mistral-7B-512k., Llama-3.1 (Dubey et al., 2024), Qwen2 (Yang et al., 2024a), Phi-3 (Abdin et al., 2024), Mistral-Nemo [^7] https://huggingface.co/mistralai/Mistral-Nemo-Instruct-2407., Jamba-1.5 (Lenz et al., 2024), Claude-3.5-Sonnet (Anthropic, 2024), Gemini-1.5 (Reid et al., 2024), and GPT-4o (Achiam et al., 2023).
[…]
Model Max Len. Recall RAG ICL Re-rank QA Summ. Cite Avg.
ProLong (8B) 512K 98.8 63.2 86.5 22.5 43.9 29.2 1.4 49.4
MegaBeam-Mistral (7B) 512K 89.6 57.0 86.2 14.7 37.3 28.9 4.0 45.4
Meta-Llama-3.1 (8B) 128K 95.2 59.5 83.9 14.0 43.2 27.0 2.9 46.5
Qwen2 (7B) 128K 38.2 45.0 77.5 3.6 36.8 6.8 2.3 30.0
Phi-3-small (7B) 128K 22.3 33.8 79.6 1.9 27.5 6.6 3.0 24.9
Mistral-Nemo (12B) 128K 14.6 40.0 84.0 0.0 22.5 18.5 0.5 25.7
Jamba-1.5-Mini (12B/52B) 256K 90.0 57.3 91.0 14.6 54.2 18.1 3.1 46.9
Meta-Llama-3.1 (70B) 128K 90.7 56.2 81.4 24.5 56.3 31.6 7.5 49.7
Claude-3.5-Sonnet 200K 94.7 38.1 61.0 7.2 12.6 36.6 18.7 38.4
Gemini-1.5-Pro 2M 91.0 71.1 79.4 59.7 59.6 46.4 43.6 64.4
GPT-4o 128K 99.9 70.2 86.3 50.0 59.3 43.2 44.3 64.8
Since most existing models do not support more than 128K tokens, to showcase ProLong’s 512K context length, we stress test ProLong on the QA and summarization tasks from 32K to 512K888 In QA and summarization, we truncate the documents at the evaluation length; hence an effective long-context model should demonstrate better performance on longer lengths.. §6.2 shows that ProLong continues to improve at a longer context window.
@@ 7 Related Work
There have been many efforts in designing more efficient architectures, for example, linear attention/RNNs (Gu & Dao, 2023; Dao & Gu, 2024; Ma et al., 2022; Sun et al., 2023; Peng et al., 2023; Yang et al., 2024b), and alternative attention architectures (Rubin & Berant, 2023; Sun et al., 2024; Yen et al., 2024). However, they often require training from scratch and many have the inherent limitations in terms of long-context recall (Jelassi et al., 2024; Arora et al., 2024). Recent works explore hybrid models (Waleffe et al., 2024; Lieber et al., 2024) or distilling existing LMs into hybrid models (Wang et al., 2024) and show promising results.
@@
Many benchmarks have been proposed for long-context evaluation (Shaham et al., 2023; Hsieh et al., 2024; Krishna et al., 2023; Zhang et al., 2024a; An et al., 2024a; Bai et al., 2024b) There are works studying particular aspects of long-context LMs as well, such as positional bias (Liu et al., 2024b), in-context learning (Bertsch et al., 2024; Li et al., 2024b), and book-length summarization (Kim et al., 2024). In this work, we follow Yen et al. (2025) for its diverse application coverage and reliable evaluations.
@@ 8 Conclusion
[…]
@@ A.1 Evaluation
Table 13: The details for our long-context evaluation following HELMET (Yen et al., 2025).
@@
Table 13 shows all the datasets we used for the long-context evaluation from HELMET (Yen et al., 2025). Note that we did not use all the datasets from HELMET for efficiency reasons and we also do not want to overfit to HELMET. We highlight some of the evaluation protocol improvements that HELMET implemented compared to previous benchmarks here:
@@
  • Model-based evaluation for long-context QA and summarization. Instead of using traditional metrics like ROUGE (which has shown to be poorly indicative of the real model performance: Deutsch & Roth, 2021; Deutsch et al., 2022; Goyal et al., 2023; Chang et al., 2024), HELMET uses model-based evaluations to compare the reference answer and the model output. For QA, HELMET uses GPT-4o to score the model output given the question and the reference answer at a 0-3 scale. For summarization, HELMET takes a similar approach as Zhang & Bansal (2021); Gao et al. (2023): it first uses GPT-4o to decompose the reference summary into atomic claims; then it uses GPT-4o to check whether each reference atomic claim is covered by the model output (recall) and whether each sentence in the model output is covered by the reference summary (precision). Yen et al. (2025) show that the model-based evaluation correlates with human perceptions significantly better than traditional metrics.
@@ A.3 Implementation details
We use various open-source packages and tools for the ProLong training and evaluation. We use PyTorch (Paszke et al., 2019) and Hugging Face transformers (Wolf et al., 2020) for the model training. We use mosaic-streaming (Mosaic ML, 2022) for loading and mixing the data and FlashAttention 2 (Dao, 2024) for efficient attention implementation. We implement sequence parallelism based on DeepSpeed-Ulysses (Jacobs et al., 2023) across groups of 8 GPUs on the same node. We only perform distributed attention if it is necessary, i.e., only on sequences of 512K length. For long-context evaluation, we use HELMET (Yen et al., 2025) and for short-context evaluation, we use lm-eval-harness (Gao et al., 2021).
@@ B.4 Instruction-tuning datasets
B.5 Synthetic data with a stronger data generator
We observe that mixing in synthetic data generated by Llama-3-8B-Instruct does not help with the long-context performance. To ensure that this is not due to the low quality of the synthetic data, we also experiment with a stronger data generator, Llama-3-70B-Instruct. We demonstrate the results in Table 23 and verify that using a stronger data generator does not change the conclusion.
Table 23: Effect of different ratios of synthetic SFT data (mixed with UltraChat). We report the 32K-and-64K-averaged performance except tasks marked with , which are evaluated at 512K for stress testing. The number of percentage is based on #tokens, not #samples. “(8B)” and “(70B)” indicate that the synthetic data are generated by Llama-3-8B-Instruct or Llama-3-70B-Instruct. Even though using synthetic data from a stronger model leads to slightly better performance than using a weaker model, only using short-context SFT data still achieves the best result in our setting.
% Synthetic Data JsonKV RAG Re-rank ICL QA Summ. Avg.
0% 65.7 58.1 38.5 80.3 49.7 42.1 55.7
1% (from 8B) 61.5 57.0 38.3 80.8 45.3 41.5 54.1
1% (from 70B) 64.7 57.3 37.4 78.4 47.0 40.8 54.2
3% (from 8B) 62.0 56.4 37.9 80.6 44.8 39.5 53.5
3% (from 70B) 65.7 57.4 38.0 80.1 48.7 42.5 55.4
10% (from 8B) 70.3 55.5 36.1 80.6 41.7 39.4 53.9
10% (from 70B) 66.3 57.0 33.4 81.2 45.3 38.4 53.6
50% (from 8B) 45.8 48.8 18.8 70.5 42.3 33.3 43.3
50% (from 70B) 55.8 53.9 23.5 74.1 50.7 39.9 49.7
B.6 Comparison to Fu et al. (2024)
We show a head-to-head comparison to the data strategy of Fu et al. (2024) in Table 24. We see that under a fair comparison, our data mix significantly outperforms Fu et al. (2024) on both short and long-context tasks. The main difference of the two data strategies is that Fu et al. (2024) proportionally up-sample long documents in each domain with an arbitrary ratio; ProLong uses a mix of short and long documents, where the ratio of the mix and the domains for the long documents are carefully ablated.
Table 24: Comparison between Fu et al. (2024) and our model. For a fair comparison, we use the same initialization (Llama-3-8B), same amount of data (5B), and same hyperparameters (§A.4). The ProLong data mix significantly outperforms Fu et al. (2024) on both short and long-context tasks.
Data Long-Context (After SFT) Short-Context (Avg.)
Recall RAG Re-rank ICL QA Summ. Avg. Before SFT After SFT
Fu et al. (2024) 95.8 52.1 23.1 72.0 31.0 37.0 51.8 64.1 65.4
Our data mix 96.0 54.9 29.4 73.9 35.7 37.9 54.6 65.5 67.5
B.7 Short-context performance after SFT
We demonstrate the detailed short-context performance of ProLong after SFT in Table 25.
Table 25: Short-context performance of our model after SFT. We also report a baseline using Llama-3-8B as the initialization and data from Fu et al. (2024), trained with 5B tokens. ProLong is initialized from Llama-3-8B-Instruct. “Llama-3-8B-Instruct + UltraChat”: for a more fair comparison to ProLong, we conduct SFT on top of Llama-3-8B-Instruct with UltraChat. ProLong largely retraines the short-context performance of Llama-3-8B-Instruct except MMLU and GSM8K. We hypothesize that the close-source instruction tuning data of Llama-3-8B-Instruct is heavily engineered to improve math and knowledge-intensive tasks, which we do not have access to. ProLong achieves comparable results to “Llama-3-8B-Instruct + UltraChat”, which further demonstrates that our data mix effective retains short-context performance.
Model HellaSwag MMLU ARC-c WinoGrande GSM8K Avg.
Llama-3-8B + Fu et al. (2024) 82.5 63.9 63.6 75.1 42.2 65.4
Llama-3-8B 82.1 66.5 59.4 77.1 44.7 66.0
Llama-3-8B-Instruct + UltraChat 82.1 65.1 64.3 75.5 60.7 69.5
ProLong 82.8 64.6 64.7 76.2 58.9 69.4
Llama-3-8B-Instruct 78.5 67.0 60.8 74.2 68.5 69.8
B.8 Evaluation on more benchmarks
[…]
Avg. MC QA Sum Diag Calc Find Number PassKey KV Avg.
ProLong (8B) 71.9 65.1 22.0 19.8 4.5 0.0 27.4 100.0 100.0 92.8 48.0
MegaBeam-Mistral 78.9 53.7 18.5 24.8 12.0 0.0 24.3 99.7 100.0 36.4 41.0
Meta-Llama-3.1 (8B) 81.3 67.2 15.5 26.7 23.0 0.0 33.1 99.5 100.0 55.0 46.7
Qwen2 26.7 39.7 5.2 15.5 8.5 0.0 24.9 76.3 94.6 0.0 29.4
Phi-3-small 72.6 71.6 8.4 24.0 20.0 0.0 31.7 100.0 100.0 19.6 41.7
Mistral-Nemo 22.7 31.9 16.8 14.3 5.5 0.0 1.4 36.6 62.7 0.0 18.8
Jamba-1.5-Mini 87.8 76.0 17.9 0.0 3.5 0.0 31.1 100.0 100.0 45.6 41.6
Meta-Llama-3.1 (70B) 75.8 75.5 23.3 31.3 18.0 0.0 43.1 99.7 100.0 2.6 43.7
GPT-4o-mini 80.8 78.2 19.1 24.8 21.5 0.0 69.7 100.0 100.0 80.4 54.9
GPT-4o 93.3 86.5 26.0 21.5 51.0 0.0 58.9 100.0 100.0 99.8 60.4
Gemini-1.5-Pro 65.3 77.5 27.7 29.0 97.5 0.0 58.0 100.0 100.0 70.4 62.2